# Data Diversity Matters for Robust Instruction Tuning

### v1

@@ Abstract
Instruction tuning has emerged as a key step in aligning large language models. One of the central challenges of instruction tuning is dataset selection, as the composition of the instruction tuning dataset can significantly impact downstream performance. In particular, researchers have hypothesized that dataset diversity and dataset quality are important indicators of downstream performance. However, it is not clear how to automatically select high quality and diverse data or how exactly quality and diversity affect instruction following ability. To resolve these issues, we propose a new algorithm, Quality-Diversity Instruction Tuning (QDIT). QDIT provides a principled algorithm to control dataset diversity and quality, allowing us to conduct an in depth study on the effect of diversity and quality on instruction tuning performance. From this study we draw two key insights (1) there is a natural tradeoff between dataset diversity and quality and (2) increasing dataset diversity significantly improves the worst case instruction following performace, therefore improving robustness. We validate the performance of QDIT on several large scale instruction tuning datasets, where we find it can improve worst case performance by 18% while maintaining or improving average performance compared to quality driven baselines.
@@ 1 Introduction
Large pre-trained language models have demonstrated a remarkable ability to perform a wide range of natural language processing tasks (Vaswani et al., 2017; Devlin et al., 2018; Radford et al., 2019; He et al., 2020; Brown et al., 2020). Although these models are powerful, pre-trained models such as GPT-3 can be quite difficult to work with and often do not follow user instructions (Brown et al., 2020). To make language models more user-friendly and to unlock instruction-following capabilities, researchers have turned to instruction tuning, in which language models are trained to follow instructions on a small set of example instruction-response pairs (Mishra et al., 2021; Wei et al., 2021; Sanh et al., 2021; Wang et al., 2022a). Instruction tuning has become extremely popular, allowing researchers to train powerful and aligned language models with a relatively small budget (Taori et al., 2023; Chiang et al., 2023; Xu et al., 2023).
One of the main challenges of instruction tuning is selecting the instruction data. Recent works have found that the composition of the instruction-tuning dataset significantly impacts the instruction following ability of the trained model, with most works identifying two qualities that an instruction tuning dataset should have: it should be high quality (Chen et al., 2023; Peng et al., 2023) and it should cover a wide range of tasks (i.e. be diverse) (Wei et al., 2021; Zhou et al., 2023; Gudibande et al., 2023). However, as of yet there has been little study of how these two factors (quality and diversity) affect on one another and how researchers can automatically select high quality and diverse datasets. Instead, current work relies on manual data curation (Zhou et al., 2023) or optimize one quantity only (Chen et al., 2023; Peng et al., 2023; Wang et al., 2022b).
@@
To bridge this gap we introduce a new algorithm, Quality-Diversity Instruction Tuning (QDIT), to select diverse and high quality instruction tuning datasets. QDIT is inspired by submodular optimization (Fujishige, 2005) and works by iteratively selecting the datapoint that will improve the joint quality-diversity score the most. QDIT provides an effective way to control the diversity and quality of the instruction tuning dataset, allowing us to conduct an in-depth study of diversity and quality in instruction tuning. From this study we identify two key findings: (1) there is an inherent tradeoff between dataset diversity and dataset quality and (2) improving dataset diversity primarily improves the worst and average case instruction following ability, while not affecting best case instruction following ability much. Based on these finding, we are able to use QDIT to optimize the quality-diversity tradeoff, improving worst case performance while maintaining or improving best case and average performance for robust instruction following. We extensively validate the QDIT algorithm with experiments on four large scale instruction tuning datasets.
@@
The rest of this article is organized as follows: Section 2 covers related literature, Section 3 presents the QDIT algorithm, Section 4 contains our experimental results, and Section 5 concludes our work.
@@ 2 Related Work
Given the widespread availability of large instruction tuning datasets (Ding et al., 2023; Zheng et al., 2023a), a new question has emerged: How should we select a training dataset from existing ones? Here data selection is necessary as training on large scale instruction tuning datasets will have a large cost and result in deterioration of model capabilities (Ouyang et al., 2022; Zhou et al., 2023; Chen et al., 2023). Several strategies have been proposed to answer this question, mainly focusing on improving dataset quality. Li et al. (2023) propose to select instructions based on difficulty. Chen et al. (2023) propose Alpagasus, in which a large language model such as ChatGPT is employed to select the highest quality (instruction, response) pairs from a large dataset. Peng et al. (2023) also seek to improve data quality, and do so by rewriting answers to instructions with GPT-4. We seek to improve upon these algorithms by including both diversity and quality in the selection criteria.
Our work is also related to several works that use submodular optimization to increase dataset diversity in natural language processing (Kumar et al., 2019; Kirchhoff and Bilmes, 2014). However, to the best of our knowledge we are the first to include both data quality and diversity in the selection procedure.
@@ 3 Method
Before presenting QDIT, we discuss how to quantify and optimize dataset diversity.
@@ 3.1 Quantifying Dataset Diversity and Quality
$d(A)=\sum_{v\in V}\textrm{max}_{a\in A}sim(a,v)$ (1)
Intuitively, we can see that a set A that has a high diversity score $d(A)$ will have an element $a\in A$ close to each element $v\in V$ and will therefore be representative of the set $V$.
In order to optimize the facility location function, researchers typically use the greedy data selection algorithm (Nemhauser et al., 1978), where the datapoint that most increases the diversity score of the current subset is selected at each iteration. The greedy algorithm is shown in Algorithm 1, where $d(a|A)=d(a\cup A)-d(A)$. We remark that this algorithm achieves the best possible approximation ratio (in the worst case) that a polynomial time algorithm can achieve unless $P=NP$ (Nemhauser et al., 1978).
To measure dataset quality, we follow prior works and measure the quality of each (instruction, response) pair using a large language model such as ChatGPT (Chen et al., 2023) or measure the quality of each datapoint using a reward model trained on large amounts of human preference data (Ouyang et al., 2022; Bai et al., 2022). We refer to such a function with $q(\cdot)$, and measure a dataset’s overall quality by averaging the quality score of each datapoint.
3.2 Quality-Diversity Instruction-Tuning
In order to select data based on both quality and diversity, we alter the greedy algorithm so that instead of selecting based on diversity alone at each iteration, we select data based on a linear combination of quality and diversity $(1-\alpha)d(a)+\alpha q(a)$. Here $\alpha$ is a hyperparameter that can be used to control the diversity tradeoff. We remark that when $\alpha=1$, QDIT becomes the quality driven selection algorithm proposed in Chen et al. (2023) and when $\alpha=0$ QDIT becomes the classical greedy algorithm (Nemhauser et al., 1978). QDIT will select points that most increase the joint diversity-quality score of the data set, therefore satisfying our goal of building a diverse and high quality dataset. The QDIT algorithm can be found in Algorithm 2. In QDIT, we use the cosine similarity of instruction embeddings as the similarity function in (1), where the instruction embeddings are computed with sentence transformers (Reimers and Gurevych, 2019). See appendix F for more details.
Algorithm 1 Greedy Data Selection
1:$K$triangleright Desired subset size
2:$A\leftarrow\emptyset$
3:$R\leftarrow V$
4:$N\leftarrow 0$
5:while $N<K$ do
6:     $a\leftarrow\textrm{argmax}_{a\in R}d(a|A)$
7:     $A\leftarrow A\cup\{a\}$
8:     $R\leftarrow R\setminus a$
9:     $N\leftarrow N+1$
10:end while
Algorithm 2 QDIT Data Selection
@@ 3.2 Quality-Diversity Instruction-Tuning
6:     $a\leftarrow\textrm{argmax}_{a\in R}(1-\alpha)*d(a|A)+\alpha*q(a)$
@@
Computational Complexity of QDIT. Finding $a\in A$ that maximizes the quality-diversity score at each time step has complexity $\mathcal{O}(|V|^{3})$, leading to a total complexity of $\mathcal{O}(|V|^{3}K)$ for QDIT based data selection. This is problematic, since for some datasets $|V|$ is greater than 1 million. We take several approaches to reduce data selection cost: (1) parallelization on GPUs, (2) employing the lazy greedy selection algorithm of Minoux (2005), and (3) sub-sampling according to Mirzasoleiman et al. (2015). These methods can be employed simultaneously, allowing us to select from millions of datapoints within a few hours. We provide more details in Appendix D, where we find that data selection has minimal cost when compared to training.
@@ 4.1 Preliminary Analysis
To qualitatively verify that the facility location function is aligned with dataset diversity, we use the Berkeley Neural Parser (Kitaev and Klein, 2018; Kitaev et al., 2019) to extract the root verb and first direct noun from each instruction in the Alpaca dataset. We then plot the distribution of verb-noun pairs in Figure 1 for random, quality driven, and QDIT data selection. From Figure 1, we observe that selecting based on the facility location function indeed improves the dataset diversity compared to random selection, while selecting based on quality alone decreases dataset diversity.
@@
Next we plot how $\alpha$ in QDIT affects dataset quality and diversity in Figure 2. Similar figures for other datasets can be found in Appendix A. From these figures, we can observe that there is a tradeoff between quality and diversity, and that QDIT allows us to smoothly control this tradeoff. Moreover, we observe that QDIT is able to improve diversity without significantly decreasing data quality. Altogether, these results indicate that QDIT is a practical way to control dataset diversity and quality.
(a) 3K Selected Points
(b) 9K Selected Points
Figure 2: Effect of $\alpha$ on dataset quality and diversity. The red line represents a randomly selected dataset.
@@ 4.2 Experimental Setup
$\diamond$ Training Setup. We use QDIT on four instruction tuning datasets: Alpaca 52K (Taori et al., 2023), Ultrachat 1.3M (Ding et al., 2023), LMSYS-Chat 1M (Zheng et al., 2023b), and a combined dataset of Alpaca 52K, Dolly 15K (Conover et al., 2023), and the OIG-small-chip2 dataset (210K). We refer to this dataset as “Mixed 270K”. For each dataset we consider two settings: a small data setting and a large data setting. For our main experiments we follow the training procedure from Taori et al. (2023) and use LLaMA-1 7B as our base model (Touvron et al., 2023). In all settings we use the same number of training epochs, meaning that the training cost is proportional to the number of training instructions. Complete hyperparameter details can be found in Appendix B.
@@
$\diamond$ Evaluating Robustness. Beyond evaluating average performance on the test dataset, we also evaluate the worst and best case performance of each model. We can evaluate the worst case performance with the HH Score by calculating the average score achieved on the worst 10% of instructions for each model (note that the worst instructions can change depending on the model). Similarly, we can evaluate the worst case performance with AlpacaFarm by measuring the losing rate on the worst 40% (smaller percents reduce the test set size by too much) of instruction categories in the WizardLM evaluation set. Here we use losing rate as a metric because a high losing rate will indicate low robustness. Best case performance is measured in a similar manner with win rate. Measuring best and worst case performance provides more detailed insights into how diversity and quality affect model robustness.
$\diamond$ Baselines. We primarily compare the QDIT algorithm with two baselines: random data selection and quality based selection. We remark that for the Alpaca dataset instruction-response quality is measured according to ChatGPT (using the same setup as Chen et al. (2023)) and for all other datasets quality is measured with the previously mentioned reward model.
We also introduce two variants of the QDIT algorithm that attempt to improve data quality and diversity. The first variant, QDIT:Threshold, first orders the data point by quality, and then iteratively selects instructions that do not have a similarity score greater than $\tau$ with any point included selected subset. This algorithm is different from QDIT in that it prevents the dataset from having similar datapoints, which QDIT allows. The second variant we propose is QDIT-Cluster, in which we first cluster the instructions and then select an equal amount of datapoints from each cluster in a quality-driven manner. Comprehensive details can be found in Appendix C, where we also provide an analysis of how these algorithms select data.
Table 1: Instruction Tuning Results. Random 50K refers to a model trained on a randomly sampled set of 50K points on the corresponding dataset. The top and bottom 40% win and lose rate is computed on the WizardLM evaluation set.
@@
Winning Score vs. Alpaca 52K normal-$\uparrow$ Winning Score vs. Random 50K normal-$\uparrow$ HH Score Mean normal-$\uparrow$ Bottom 10% HH Score normal-$\uparrow$ Bottom 40% Lose Rate normal-$\downarrow$ Top 10% HH Scorenormal-$\uparrow$ Top 40% Win Rate normal-$\uparrow$
Random 9K 0.975 - 3.756 0.489 19.23 6.942 21.77
Quality 9K 1.013 - 3.944 0.59 30.38 7.478 31.82
QDIT 9K 1.013 - 3.922 0.573 29.55 7.437 25.71
Random 3K 0.928 - 3.601 0.368 37.35 6.876 26.25
Quality 3K 0.923 - 3.732 0.352 30.63 6.972 20.45
@@
QDIT:Threshold 3K 0.986 - 3.745 0.452 25.99 6.943 29
QDIT 3K 0.961 - 3.809 0.5 25 7.03 27.06
Winning Score vs. Alpaca 52K normal-$\uparrow$ Winning Score vs. Random 50K normal-$\uparrow$ HH Score Mean normal-$\uparrow$ Bottom 10% HH Score normal-$\uparrow$ Bottom 40% Lose Rate normal-$\downarrow$ Top 10% HH Scorenormal-$\uparrow$ Top 40% Win Rate normal-$\uparrow$
Random 50K 0.927 1 3.6 0.396 29.06 7.002 25.89
Quality 50K 0.973 1.065 4.295 1.186 34.78 7.746 28.7
QDIT 50K 0.975 1.051 4.33 1.223 35.22 7.761 26.32
Random 10K 0.896 0.985 3.609 0.314 32.69 6.863 24.1
Quality 10K 0.956 1.042 4.342 1.117 33.33 7.952 30
@@
QDIT:Threshold 10K 0.98 1.06 4.282 1.051 33.96 8.005 30.7
QDIT 10K 0.982 1.063 4.429 1.3 32.65 8.006 26.74
Winning Score vs. Alpaca 52K normal-$\uparrow$ Winning Score vs. Random 50K normal-$\uparrow$ HH Score Mean normal-$\uparrow$ Bottom 10% HH Score normal-$\uparrow$ Bottom 40% Lose Rate normal-$\downarrow$ Top 10% HH Scorenormal-$\uparrow$ Top 40% Win Rate normal-$\uparrow$
Random 50K 1.189 1 4.866 0.811 25 8.594 50
Quality 50K 1.249 1.068 5.91 2.116 21.65 9.562 56.25
QDIT 50K 1.26 1.094 5.909 2.185 22.52 9.69 57.45
Random 10K 1.137 0.968 4.891 0.931 27.47 8.756 47.78
Quality 10K 1.228 1.011 5.844 1.846 25.47 9.74 48.15

### v3

@@ Abstract
Recent works have shown that by curating high quality and diverse instruction tuning datasets, we can significantly improve instruction-following capabilities. However, creating such datasets is difficult and most works rely on manual curation or proprietary language models. Automatic data curation is difficult as it is still not clear how we can define diversity for instruction tuning, how diversity and quality depend on one other, and how we can optimize dataset quality and diversity. To resolve these issue, we propose a new algorithm, Quality-Diversity Instruction Tuning (QDIT). QDIT provides a simple method to simultaneously control dataset diversity and quality, allowing us to conduct an in-depth study on the effect of diversity and quality on instruction tuning performance. From this study we draw two key insights (1) there is a natural tradeoff between data diversity and quality and (2) increasing data diversity significantly improves the worst case instruction following performance, therefore improving robustness. We validate the performance of QDIT on several large scale instruction tuning datasets, where we find it can substantially improve worst and average case performance compared to quality-driven data selection.
@@ 1 Introduction
Large pre-trained language models have demonstrated a remarkable ability to perform a wide range of natural language processing tasks (Vaswani et al., 2017; Devlin et al., 2018; Radford et al., 2019; He et al., 2020; Brown et al., 2020). Although these models are powerful, pre-trained models such as GPT-3 can be quite difficult to work with and often do not follow user instructions (Brown et al., 2020). To unlock instruction-following capabilities, researchers have turned to instruction tuning, in which language models are trained to follow instructions on a small set of example instruction-response pairs (Mishra et al., 2021; Wei et al., 2021; Sanh et al., 2021; Wang et al., 2022a). Instruction tuning has become extremely popular, as it provides a simple way for researchers to train powerful and aligned language models (Taori et al., 2023; Chiang et al., 2023; Xu et al., 2023).
Although initial works apply instruction tuning to large-scale datasets, it has recently been found that a small set of well chosen instruction-response pairs is sufficient for good performance. In particular, Zhou et al. (2023) showed that by training on only 1000 instructions manually selected or crafted by experts, superior performance can be achieved compared to training on larger datasets. Training with a small dataset has the added benefits of lowering training costs and enabling faster iteration. Although such manual data selection is not scalable, this work raises an important question: How can we automatically select an instruction tuning dataset?
Recent work on dataset curation have identified two characteristics that an instruction tuning dataset should have: (1) the instruction responses should be high quality (Chen et al., 2023; Peng et al., 2023) and (2) the instructions should cover a wide range of tasks (i.e. be diverse) (Wei et al., 2021; Zhou et al., 2023; Gudibande et al., 2023). To curate high quality datasets, researchers have used proprietary LLMs to measure the quality of each data point in the dataset and then select only the highest quality data points. To improve dataset diversity, researchers have manually selected instructions that cover a wide range of topics and formats (Zhou et al., 2023; Ivison et al., 2023). While both of these methods enhance instruction-following capabilities, it is not clear how we can select high quality and diverse datasets without relying on manual curation from human experts, a process that is time-consuming and expensive.
In pursuit of this goal, we propose a new algorithm, QDIT, to measure and optimize the diversity and quality of instruction tuning datasets. QDIT measures diversity using the facility location function (Cornuéjols et al., 1983). The facility location function provides an intuitive measure of subset diversity, as it essentially measures how well represented each data point in the full dataset is by the data points in the selected subset. With this diversity function and quality functions from prior works, we then define a dataset’s quality-diversity score as a simple linear combination of dataset quality and diversity. To optimize the quality-diversity score, QDIT employs a greedy strategy, where the data point that will improve the joint quality-diversity score the most is selected at each time step (Nemhauser et al., 1978). This procedure is extremely efficient, and can easily scale to datasets with millions of instruction.
@@
QDIT provides an effective way to control the diversity and quality of the instruction tuning dataset, allowing us to conduct an in-depth study of diversity and quality in instruction tuning. From this study we identify two key findings: (1) there is an inherent tradeoff between dataset diversity and dataset quality and (2) improving dataset diversity primarily improves the worst and average case instruction following ability, while not affecting best case instruction following ability much. Based on these finding, we are able to use QDIT to optimize the quality-diversity tradeoff, improving worst case performance while maintaining or improving best case and average performance for robust instruction following. We extensively validate our results on five large-scale instruction tuning datasets.
@@
The rest of this paper is organized as follows: Section 2 covers related literature, Section 3 presents the QDIT algorithm, Section 4 contains our experimental results, and Section 5 concludes our work.
@@ 2 Related Work
There have been several works that attempt to improve the quality and diversity of instruction tuning datasets.
$\diamond$ Manual Data Selection. Several works have shown that superior instruction-following capabilities can be unlocked by carefully selecting and writing instruction-response pairs (Zhou et al., 2023; Touvron et al., 2023; Wang et al., 2023a; Ivison et al., 2023). To select such data quality and diversity are emphasized, with data from various scientific fields and internet forums being selected. However it is not clear how such datasets can be automatically selected.
$\diamond$ Distilling Closed Models. To reduce the human effort required for dataset creation, researchers have used powerful proprietary LLMs such as GPT-4 to create instruction tuning datasets (Taori et al., 2023; Peng et al., 2023; Chia et al., 2023). Again researchers found dataset quality (Peng et al., 2023) and diversity (Xu et al., 2023; Li et al., 2023a) to be most important. Although resulting in powerful datasets, the reliance on proprietary language models in these works is expensive and may raise legal concerns (Wang et al., 2023b).
$\diamond$ Automatic Data Selection for Instruction Tuning. Due to the aforementioned issues, in this paper we focus on automatic selection of smaller instruction tuning datasets from larger ones. Chen et al. (2023); Dong et al. (2023a) show that by rating the quality of each data point and training on the highest quality data points, downstream performance can be significantly improved. Li et al. (2023b) propose to select instructions based on difficulty. A few works attempt to increase diversity by restricting the distance between selected points to be larger than a given threshold (Wang et al., 2022b; Liu et al., 2023). We find that this method does not necessarily lead to a significant increase in diversity, and that they are more similar to data de-duplication (Tirumala et al., 2023). We compare QDIT to similar approaches in our experiments.
Our work is also related to several works that use submodular optimization to increase dataset diversity in natural language processing (Kumar et al., 2019; Kirchhoff and Bilmes, 2014). To the best of our knowledge, we are the first to use both data quality and diversity in the selection procedure.
@@ 3 Method
Before presenting QDIT, we discuss how to quantify instruction diversity and instruction-response quality.
@@ 3.1 Quantifying Dataset Diversity and Quality
$d(A)=\sum_{v\in V}\textrm{max}_{a\in A}\mathrm{sim}(a,v),$ (1)
where $sim(a,v)$ refers to the similarity of $a$ and $v$. In QDIT, we use the cosine similarity of instruction embeddings as the similarity function in (1), where the instruction embeddings are computed with sentence transformers (Reimers and Gurevych, 2019). See Appendix A for more details. Intuitively, we can see that a set $A$ that has a high diversity score $d(A)$ will have an $a\in A$ close to each $v\in V$ and will therefore be representative of the set $V$.
To measure dataset quality, we follow prior works and measure the quality of each (instruction, response) pair using a large language model such as ChatGPT (Chen et al., 2023) or measure the quality of each data point using a scoring model trained on large amounts of human preference data (Ouyang et al., 2022; Bai et al., 2022). We refer to such a function with $q(\cdot)$, and measure a dataset’s overall quality by averaging the quality score of each data point.
3.2 Quality-Diversity Instruction Tuning
In order to simultaneously control quality and diversity of the selected data, we propose a linear combination of quality and diversity as the quality-diversity (Q-D) score:
$f(a|A,\alpha)=(1-\alpha)d(a|A)+\alpha q(a),$
where $\alpha\in[0,1]$ is a hyperparameter controlling the tradeoff betwen quality and diversity.
To optimize the Q-D score of the selected data, we consider a greedy algorithm – named QDIT shown in Algorithm 1. Specifically, at each iteration, we select the data point that most increases the Q-D score of the current subset. When $\alpha=0$, the greedy algorithm achieves the best possible approximation ratio (in the worst case) that a polynomial time algorithm can achieve. See more details in (Nemhauser et al., 1978). We remark that when $\alpha=1$, QDIT is reduced to the quality driven selection algorithm proposed in Chen et al. (2023) and when $\alpha=0$, QDIT is reduced the classical greedy algorithm (Nemhauser et al., 1978).
Algorithm 1 QDIT Data Selection: Select a subset of $K$ data points from $N$ data points
@@ 3.2 Quality-Diversity Instruction Tuning
6:     $a\leftarrow\textrm{argmax}_{a\in R}f(a|A,\alpha)$
@@
Once we finish the selection, we apply instruction tuning on the selected data.
Computational Complexity of QDIT. Finding $a\in A$ that maximizes the quality-diversity score at each time step has complexity $\mathcal{O}(|V|^{3})$, leading to a total complexity of $\mathcal{O}(|V|^{3}K)$ for QDIT based data selection. This is problematic, since for some datasets $|V|$ is greater than 1 million. We take several approaches to reduce data selection cost: (1) parallelization on GPUs, (2) employing the lazy greedy selection algorithm of Minoux (2005), and (3) sub-sampling according to Mirzasoleiman et al. (2015). These methods can be employed simultaneously, allowing us to select from millions of data points within a few hours. We provide more details in Appendix B, where we find that data selection has minimal cost compared to training.
@@ 4.1 Preliminary Analysis
To qualitatively verify that the facility location function is aligned with dataset diversity, we use the Berkeley Neural Parser (Kitaev and Klein, 2018; Kitaev et al., 2019) to extract the root verb and first direct noun from each instruction in the Alpaca dataset. We then plot the distribution of verb-noun pairs in Figure 1 for random, quality driven, and QDIT data selection. From Figure 1, we observe that selecting based on the facility location function indeed improves the dataset diversity compared to random selection, as more verb-noun pairs are included in the dataset and the dataset becomes more uniform. On the other hand, selecting based on quality alone decreases dataset diversity.
@@
Next we plot how $\alpha$ in QDIT affects dataset quality and diversity of 3K selected points in Figure 2. Similar figures for other datasets can be found in Appendix C. From these figures, we can observe that there is a tradeoff between quality and diversity, and that QDIT allows us to smoothly control this tradeoff. Moreover, we observe that QDIT is able to improve diversity without significantly decreasing data quality. Altogether, these results indicate that QDIT is a practical way to control dataset diversity and quality.
Figure 2: Effect of $\alpha$ on QDIT’s dataset quality and diversity. The red line represents a randomly selected dataset.
@@ 4.2 Experimental Setup
$\diamond$ Training Setup. We use QDIT on two small instruction tuning datasets: Alpaca 52K (Taori et al., 2023), and Dolly 15K (Conover et al., 2023) as well as three large scale datasets: Ultrachat 1.3M (Ding et al., 2023), LMSYS-Chat 1M (Zheng et al., 2023), and a combined dataset of Alpaca 52K, Dolly 15K , and the OIG-small-chip2 dataset (210K). We refer to this dataset as “Mixed 270K”. For each large dataset we select a dataset size of 10K points. For each small dataset size we follow the small data setting from Chen et al. (2023) and select $\sim 5\%$ of the original dataset.
To measure instruction-response quality, we use the provided ChatGPT quality scores from Chen et al. (2023) for Alpaca and for all other datasets we use the reward model from Dong et al. (2023b), which is trained on the Anthropic Helpful Harmless dataset and achieves a test accuracy of over 75%. For our main experiments we follow the training procedure from Taori et al. (2023) and use LLaMA-1 7B as our base model (Touvron et al., 2023). In all settings we use the same number of training epochs, meaning that the training cost is proportional to the number of training instructions. Complete hyperparameter details can be found in Appendix D.
@@
$\diamond$ Evaluating Robustness. Beyond evaluating average performance on the test dataset, we also evaluate the worst and best case performance of each model. We can evaluate the worst case performance with the HH Score by calculating the average score achieved on the worst 10% of instructions for each model (note that the worst instructions can change depending on the model). Similarly, we can evaluate the worst case performance with pairwise comparison by measuring the winning score on the hardest 10% of prompts according to HH score. Best case performance is measured in a similar manner. Measuring best and worst case performance provides more detailed insights into how diversity and quality affect model robustness.
$\diamond$ Baselines. We primarily compare the QDIT algorithm with two baselines: random data selection and quality based selection. For the Alpaca dataset, the quality baseline is trained on the same data as Alpagasus (Chen et al., 2023).
Table 1: Instruction Tuning results. Random 50K refers to a model trained on a randomly sampled set of 50K points on the corresponding dataset. The top and bottom 10% winning and losing score is computed versus the Alpaca 52K.
[…]
Random 10K 1.138 0.969 6.219 2.620 1.074 9.526 1.167
Quality 10K 1.224 1.013 6.961 3.405 1.175 10.454 1.303
QDIT 10K 1.226 1.038 6.993 3.497 1.280 10.454 1.293
[…]
Random 10K 0.899 0.986 5.443 2.260 0.940 8.80 0.896
Quality 10K 0.959 1.04 6.140 2.973 0.989 9.670 0.984
QDIT 10K 0.987 1.083 6.276 3.054 0.973 9.676 1.044
[…]
Random 10K 1.113 1.0 6.176 2.575 1.057 9.589 1.187
Quality 10K 1.198 1.137 7.066 3.391 1.052 10.39 1.284
QDIT 10K 1.224 1.149 7.0 3.551 1.149 10.34 1.343
Table 2: Instruction Tuning results on small datasets. The top and bottom 10% winning and losing score is computed versus the Alpaca 52K.
@@
Random 3K 0.929 - 5.486 2.105 0.931 8.986 0.937
Quality 3K 0.920 - 5.629 2.306 0.873 9.073 0.934
QDIT 3K 1.026 - 5.661 2.513 0.924 8.791 1.056
Dolly 15K Average Performance Worst Case Performance Best Case Performance
@@
Random 1K 0.64 0.72 4.632 1.474 0.632 6.144 0.645
Quality 1K 0.71 0.83 5.389 2.082 0.681 7.751 0.746
QDIT 1K 0.739 0.874 5.495 2.229 0.742 7.873 0.872
4.3 Main Results
The main experimental results can be found in Table 1 for large datasets and Table 2 for small datasets.
Effect of Quality. Similar to prior works, we find that selecting data based on quality significantly improves average performance, improving the average winning score versus Alpaca 52K by 6.37% and the average HH score by 11.6% when compared to random selection. However, we also find that selecting based on quality alone can hurt worst case performance, decreasing the lowest 10% winning score in two out of five settings compared to random data selection. We hypothesize that this performance drop is due to the fact that quality based selection hurts dataset diversity.
Figure 3: Percent improvement in HH Score of QDIT over Quality-based selection. Performance is averaged over the five datasets and the shaded area represents one standard deviation.
Effect of Diversity. On the other hand, we find that data selection with QDIT is able to achieve both a high average HH score (QDIT improves upon quality driven selection by 4.17% for winning score vs Alpaca 52K and improves average HH score by 1.5%) while achieving a much better worst case performance than both random and quality-driven selection. In particular, we find that QDIT improves worst case HH score by 5.2% and worst winning score vs Alpaca 52K by 6.26% when compared to quality-driven data selection. This trend can be seen in Figure 3, where QDIT improves most over quality selection in the lowest and middle percentiles, while not affecting the highest percentiles. We hypothesize that QDIT’s more diverse dataset teaches the model to respond to a wide range of instructions, thereby decreasing the probability that it fails to follow evaluation instructions.
From these experiments we conclude that by increasing data diversity while maintaining data quality, QDIT can improve instruction following capability compared to quality-driven selection. We remark that improvements in worst-case performance are typically more important than improvements in best-case performance, as a high worst case performance will ensure users have a consistently positive experience.
(a) Dolly 1K
(b) Alpaca 3K
(c) Mixed 10K
(d) Ultrachat 10K
(e) LMSYS 10K
Figure 4: Effect of $\alpha$ on best case and worst case performance. The red line represents a randomly selected dataset and $\alpha=1.0$ is quality-driven data selection. Worst HH Score refers to the bottom 10 percent of HH scores.
Table 3: Evaluation on benchmark datasets. We bold the best result out of quality based selection and QDIT based selection.
Ultrachat 10K LMSYS 10K
MMLU BBH ARC DROP LAMBADA SCIQ AVG MMLU BBH ARC DROP LAMBADA SCIQ AVG
Random 32.12 33.19 58.34 26.24 69.77 85.4 50.84 33.05 32.57 60.15 25.06 68.5 86.7 51.01